### v3

@@ D.1 The Case of $\mu_{f}>0$
$=2+\frac{1}{\eta+2\kappa_{f}-1}\lor\frac{1}{0.5+2\kappa_{f}}\leq 4+\frac{1}{\eta+2\kappa_{f}-1},$
$\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\begin{cases}\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{t-1}&t\leq\tau\\ \left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(t-\tau+1+4\kappa_{f})(t-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
Thus, there is
$F(x^{T+1})-F(x)\leq\left(8+\frac{2}{\eta+2\kappa_{f}-1}\right)\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (45), we can obtain
@@ D.2 The Case of $\mu_{h}>0$
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
@@
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.2 holds, we will have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}$
$=\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$ (48)
Observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\eta_{t}\Gamma_{t-1}=\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{h}},$
where $\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}\coloneqq\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})$. So (48) can be rewritten as
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}.$ (49)
We can check that
@@
Therefore, (49) is true for all cases.
If $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, we have
$\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}=\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})=\frac{(t+1+4\kappa_{h})(t+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}.$
Hence, by (49),
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\frac{(T+1+4\kappa_{h})(T+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}+\frac{8(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{\frac{1}{(t+4\kappa_{h})^{2}}}{\frac{(T+1+4\kappa_{h})(T+2+4\kappa_{h})}{(t+4\kappa_{h})(t+1+4\kappa_{h})}-1}$
$=$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{t+1+4\kappa_{h}}{t+4\kappa_{h}}\cdot\frac{1}{(T+1-t)(T+2+8\kappa_{h}+t)}$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{1}{2T+3+8\kappa_{h}}\left(\frac{1}{T+1-t}+\frac{1}{T+2+8\kappa_{h}+t}\right)$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}}\cdot\frac{1+\log T+\log\frac{2T+2+8\kappa_{h}}{T+2+8\kappa_{h}}}{2T+3+8\kappa_{h}}$
$\leq$ $\frac{(1+4\kappa_{h})(2+4\kappa_{h})\mu_{h}D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{h}(2T+3+8\kappa_{h})}$
@@
If $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$, we know

### v4

@@ D.1 The Case of $\mu_{f}>0$
$=2+\frac{1}{\eta+2\kappa_{f}-1}\lor\frac{1}{0.5+2\kappa_{f}}\leq 4+\frac{1}{\eta+2\kappa_{f}-1}.$
Thus, using (44), we obtain
$F(x^{T+1})-F(x)\leq\left(8+\frac{2}{\eta+2\kappa_{f}-1}\right)\left(\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (43), we can conclude
@@
D.1.1 Optimal Rates via a New Step Size
Now, we present the full version of Theorems 3.8 and 3.9.
Theorem D.3 (Full version of Theorem 3.8).
Under Assumptions 2.-4. and 5A. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$, for any $x\in\mathcal{X}$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+\kappa_{f})}&t\geq\tau_{2}+1\end{cases}$ with $\tau_{1}\coloneqq\left\lceil\frac{T}{4}\right\rceil$ and $\tau_{2}\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{4+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}}{\mu_{f}(T+\kappa_{f})}\right).$
Proof.
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (45)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}$. Now, let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$:
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}=\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+\kappa_{f})}\leq\frac{1}{2L}\land\frac{1}{\mu(T+\kappa_{f})}\leq\frac{1}{2L\lor\mu_{f}}&t\geq\tau_{2}+1\end{cases}.$
Therefore, (45) holds.
In the following proof, we write
$\eta=1$ and $\eta_{*}=\frac{1}{2L(T-\tau_{2})}\land\frac{1}{\mu_{f}(T-\tau_{2})(T+\kappa_{f})}.$ (46)
Note that we have $\eta_{1}=\frac{1}{\mu_{f}(\eta+2\kappa_{f})}$ and
$\gamma_{t\in\left[T\right]}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}=\begin{cases}\frac{\eta_{1}}{(1-\mu_{f}\eta_{1})^{t-1}}&t\leq\tau_{1}\\ \frac{(t-\tau_{1}+1+4\kappa_{f})}{(1-\mu_{f}\eta_{1})^{\tau_{1}-1}\mu_{f}(1+2\kappa_{f})(1+4\kappa_{f})}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{(\tau_{2}-\tau_{1}+1+4\kappa_{f})(\tau_{2}-\tau_{1}+2+4\kappa_{f})\eta_{*}(T-t+1)}{(1-\mu_{f}\eta_{1})^{\tau_{1}-1}(1+4\kappa_{f})(2+4\kappa_{f})\prod_{s=\tau_{2}+1}^{t}(1-\mu_{f}\eta_{s})}&t\geq\tau_{2}+1\end{cases}.$ (47)
So, there is
[…]
$\leq\frac{(1-\mu_{f}\eta_{1})^{\tau_{1}}}{\eta_{1}}\leq(\eta_{1}^{-1}-\mu_{f})\exp\left(-\mu_{f}\eta_{1}(\tau_{1}-1)\right)$
$\overset{\eta_{1}\mu_{f}\leq 1,\tau_{1}\geq\frac{T}{4}}{\leq}e(\eta_{1}^{-1}-\mu_{f})\exp\left(-\frac{\mu_{f}\eta_{1}T}{4}\right)=2eL\exp\left(-\frac{T}{4+8\kappa_{f}}\right).$ (48)
Now, we observe that
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}$ $=\sum_{t=1}^{\tau_{1}}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}+\sum_{t=\tau_{1}+1}^{\tau_{2}}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}+\sum_{t=\tau_{2}+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}$
$=\underbrace{\sum_{t=1}^{\tau_{1}}\frac{\gamma_{t}\eta_{1}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{I}}+\underbrace{\sum_{t=\tau_{1}+1}^{\tau_{2}}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{II}}+\underbrace{\sum_{t=\tau_{2}+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{III}}.$
We bound terms $\mathrm{I}$, $\mathrm{II}$, and $\mathrm{III}$ separately in the following.
  • Term $\mathrm{I}$: When $t\leq\tau_{1}$, we note that
[…]
    $=\frac{2\mu_{f}(1+2\kappa_{f})(1+4\kappa_{f})\eta_{1}(1-\mu_{f}\eta_{1})^{\tau_{1}-t}}{(\tau_{2}-\tau_{1})(\tau_{2}-\tau_{1}+3+8\kappa_{f})},$
    implying that
    $\mathrm{I}$ $\leq\frac{2\mu_{f}(1+2\kappa_{f})(1+4\kappa_{f})\eta_{1}^{2}\sum_{t=1}^{\tau_{1}}(1-\mu_{f}\eta_{1})^{\tau_{1}-t}}{(\tau_{2}-\tau_{1})(\tau_{2}-\tau_{1}+3+8\kappa_{f})}\leq\frac{2\mu_{f}(1+2\kappa_{f})(1+4\kappa_{f})\eta_{1}^{2}(\tau_{1}\land\frac{1}{\mu_{f}\eta_{1}})}{(\tau_{2}-\tau_{1})(\tau_{2}-\tau_{1}+3+8\kappa_{f})}$
[…]
    $\overset{\tau_{1}\leq\frac{T+3}{4},\tau_{2}\geq\frac{T}{2},T\geq 4}{\leq}\frac{112}{\mu_{f}(T+9+32\kappa_{f})}\leq\frac{112}{\mu_{f}(T+\kappa_{f})}.$
  • Term $\mathrm{II}$: When $\tau_{1}+1\leq t\leq\tau_{2}$, we note that
[…]
    $\leq\frac{2(t-\tau_{1}+1+4\kappa_{f})}{\mu_{f}(\tau_{2}-\tau_{1}+1+4\kappa_{f})(\tau_{2}-\tau_{1}+2+4\kappa_{f})\eta_{*}\sum_{s=\tau_{2}+1}^{T}T-s+1}$
    $=\frac{2(t-\tau_{1}+1+4\kappa_{f})}{\mu_{f}(\tau_{2}-\tau_{1}+1+4\kappa_{f})(\tau_{2}-\tau_{1}+2+4\kappa_{f})\eta_{*}(T-\tau_{2})(T-\tau_{2}+1)},$
    implying that
    $\mathrm{II}$ $\leq\frac{4(\tau_{2}-\tau_{1})}{\mu_{f}^{2}(\tau_{2}-\tau_{1}+1+4\kappa_{f})(\tau_{2}-\tau_{1}+2+4\kappa_{f})\eta_{*}(T-\tau_{2})(T-\tau_{2}+1)}$
[…]
    $\overset{(a)}{\leq}\frac{3(2\kappa_{f}\lor(T+\kappa_{f}))}{\mu_{f}(\tau_{2}-\tau_{1}+1+4\kappa_{f})(\tau_{2}-\tau_{1}+2+4\kappa_{f})}\overset{(b)}{\leq}\frac{48(2\kappa_{f}\lor(T+\kappa_{f}))}{\mu_{f}(T+1+16\kappa_{f})(T+5+16\kappa_{f})}$
    $\leq\frac{54}{\mu_{f}(T+\kappa_{f})},$
    where $(a)$ is due to $\frac{\tau_{2}-\tau_{1}}{T-\tau_{2}+1}\leq\frac{\frac{T+1}{2}-\frac{T}{4}}{\frac{T+1}{2}}=\frac{T+2}{2(T+1)}\leq\frac{3}{4}$, and $(b)$ is by $\tau_{2}-\tau_{1}\geq\frac{T}{2}-\frac{T+3}{4}=\frac{T-3}{4}$.
  • Term $\mathrm{III}$: When $t\geq\tau_{2}+1$, we note that $\frac{\gamma_{t}}{\eta_{t}}=\frac{1}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}$ is always non-decreasing. Hence, we have
    $\frac{\gamma_{t}}{\sum_{s=t}^{T}\gamma_{s}}=\frac{\frac{\gamma_{t}}{\eta_{t}}\eta_{t}}{\sum_{s=t}^{T}\frac{\gamma_{s}}{\eta_{s}}\eta_{s}}\leq\frac{\eta_{t}}{\sum_{s=t}^{T}\eta_{s}}=\frac{T-t+1}{\sum_{s=t}^{T}T-s+1}=\frac{2}{T-t+2},$
    implying that
[…]
Therefore, we obtain
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\leq\frac{112}{\mu_{f}(T+\kappa_{f})}+\frac{54}{\mu_{f}(T+\kappa_{f})}+\frac{2}{\mu_{f}(T+\kappa_{f})}=\frac{168}{\mu_{f}(T+\kappa_{f})}.$ (49)
We combine (45), (48), and (49) to conclude
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{2eLD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{4+8\kappa_{f}}\right)}+\frac{336(M^{2}+\sigma^{2})}{\mu_{f}(T+\kappa_{f})}=O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{4+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}}{\mu_{f}(T+\kappa_{f})}\right).$ (50)
Theorem D.4 (Full version of Theorem 3.9).
Under Assumptions 2.-4. and 5B. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$ and $\delta\in(0,1)$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(\eta+2\kappa_{f})}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+2\kappa_{f})}&t\geq\tau_{2}+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{f}>1$, $\tau_{1}\coloneqq\left\lceil\frac{T}{4}\right\rceil$ and $\tau_{2}\coloneqq\left\lceil\frac{T}{2}\right\rceil$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x)\leq O\left(\left(1\lor\frac{1}{\eta+2\kappa_{f}-1}\right)\left[\frac{\mu_{f}(\eta+\kappa_{f})D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{4\eta+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}\log\frac{1}{\delta}}{\mu_{f}(T+\kappa_{f})}\right]\right).$
Proof.
When $\mu_{f}>0$ and $\mu_{h}=0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ in Lemma 4.3 holds, we have, with probability at least $1-\delta$,
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}$. Now, let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$:
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(\eta+2\kappa_{f})}\leq\frac{1}{\mu_{f}(2\kappa_{f}\lor(\eta+\kappa_{f}))}\leq\frac{1}{2L\lor\mu_{f}}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+2\kappa_{f})}\leq\frac{1}{2L}\land\frac{1}{\mu(T+2\kappa_{f})}\leq\frac{1}{2L\lor\mu_{f}}&t\geq\tau_{2}+1\end{cases}.$
Therefore, (51) holds.
Note that we have
$\eta_{\tau_{1}+1}=\frac{1}{\mu_{f}(1.5+2\kappa_{f})}\geq\frac{1}{\mu_{f}(T+2\kappa_{f})}\geq\eta_{\tau_{2}+1},$
which implies that
$1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}$ $=1+\frac{1}{1-\mu_{f}(\eta_{2}\lor\eta_{\tau_{1}+1}\lor\eta_{\tau_{2}+1})}=1+\frac{1}{1-\mu_{f}(\eta_{2}\lor\eta_{\tau_{1}+1})}$
$=2+\frac{1}{\eta\land 1.5+2\kappa_{f}-1}=2+\frac{1}{\eta+2\kappa_{f}-1}\lor\frac{1}{0.5+2\kappa_{f}}$
$\leq 4+\frac{1}{\eta+2\kappa_{f}-1}.$
Thus, using (51), we obtain
$F(x^{T+1})-F(x)\leq\left(8+\frac{2}{\eta+2\kappa_{f}-1}\right)\left(\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
Following similar steps in the proof of (50), we can conclude
$F(x^{T+1})-F(x)\leq O\left(\left(1\lor\frac{1}{\eta+2\kappa_{f}-1}\right)\left[\frac{\mu_{f}(\eta+\kappa_{f})D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{4\eta+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}\log\frac{1}{\delta}}{\mu_{f}(T+\kappa_{f})}\right]\right).$
@@ D.2 The Case of $\mu_{h}>0$
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}(1+\kappa_{h})^{2}D_{\psi}(x,x^{1})}{T(T+\kappa_{h})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(\frac{1}{\eta+2\kappa_{h}}\lor\log T\right)\frac{M^{2}+\sigma^{2}}{\mu_{h}(T+\kappa_{h})}\right).$
@@
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.2 holds, we have
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}$
$=\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (52)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})$. Now, let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L}$ for the two choices respectively:
@@
Therefore, (52) holds for both cases.
If $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{h}(t+4\kappa_{h})}$, we have $\eta_{1}=\frac{2}{\mu_{h}(1+4\kappa_{h})}$ and
$\gamma_{t\in\left[T\right]}$ $=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\frac{2}{\mu_{h}(t+4\kappa_{h})}\prod_{s=2}^{t}\left(1+\frac{2}{s-1+4\kappa_{h}}\right)$
$=\frac{2(t+1+4\kappa_{h})}{\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})}=\frac{t+1+4\kappa_{h}}{2+4\kappa_{h}}\eta_{1}.$ (53)
Hence, by (52) and (53),
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(2+4\kappa_{h})\eta_{1}^{-1}D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}t+1+4\kappa_{h}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{(t+1+4\kappa_{h})\eta_{t}}{\sum_{s=t}^{T}s+1+4\kappa_{h}}$
$=\frac{\mu_{h}(2+4\kappa_{h})(1+4\kappa_{h})D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{t+1+4\kappa_{h}}{(t+4\kappa_{h})(T-t+1)(T+t+2+8\kappa_{h})}$
$\leq\frac{\mu_{h}(2+4\kappa_{h})(1+4\kappa_{h})D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}}\sum_{t=1}^{T}\frac{1}{(T-t+1)(T+t+2+8\kappa_{h})}$
$=\frac{\mu_{h}(2+4\kappa_{h})(1+4\kappa_{h})D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}(2T+3+8\kappa_{h})}\sum_{t=1}^{T}\frac{1}{T-t+1}+\frac{1}{T+t+2+8\kappa_{h}}$
$\leq\frac{\mu_{h}(2+4\kappa_{h})(1+4\kappa_{h})D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})}{\mu_{h}(2T+3+8\kappa_{h})}\left(1+\log T+\log\frac{2T+2+8\kappa_{h}}{T+2+8\kappa_{h}}\right)$
$\leq\frac{\mu_{h}(2+4\kappa_{h})(1+4\kappa_{h})D_{\psi}(x,x^{1})}{T(T+3+8\kappa_{h})}+\frac{16(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{h}(2T+3+8\kappa_{h})}$
@@
If $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{h}(\eta+2\kappa_{h})}&t\leq\tau\\ \frac{2}{\mu_{h}(t-\tau+4\kappa_{h})}&t\geq\tau+1\end{cases}$, where $\eta\geq 0$ satisfies $\eta+\kappa_{h}>0$ and $\tau=\left\lceil\frac{T}{2}\right\rceil$, we know $\eta_{1}=\frac{1}{\mu_{h}(\eta+2\kappa_{h})}$ and